How do you solve i^199 algebraically. PLease show all the math
Answer by Alan3354(69443)

i^4 = 1
i^199 = i^3 * i^196 = i^3 * (i^4)^49 = i^3 * 1^49
i^199 = i^3 = i * i^2 = -i
--------------
To sum it up:
Divide the exponent by 4 and keep the remainder
199/4 = 49 + remainder 3
= i^3